# Recovery of rational functions via Hankel pencil method and sensitivities of the poles

### v2

@@ 3.1 Sensitivity of Hankel pencil eigenvalues with unstructured perturbations
3.2 Sensitivity Hankel pencil eigenvalues with structured perturbations
In this section we consider structured perturbations of the GEP (3.1) for the matrix pencil (2.6), i.e. the perturbation matrices have a particular structure, they are again Hankel matrices. We give a short description of the results obtained in [24]. Let $\boldsymbol{\delta}_{2M}:=(\delta_{0},\delta_{1},\ldots,\delta_{2M-1})^{T}$ be a vector of perturbations and we define
$\Delta_{\mathbf{H}_{M}(0)}:=(\delta_{k+\ell})_{k,\ell=0}^{M-1,M-1},\ \Delta_{\mathbf{H}_{M}(1)}:=(\delta_{k+\ell+1})_{k,\ell=0}^{M-1,M-1}.$
@@ 3.2 Sensitivity Hankel pencil eigenvalues with structured perturbations
taking into account (3.12), we define vectors $\boldsymbol{\mathcal{S}}^{(j)}_{2M}:=\left(S^{(j)}_{1},\ldots,S^{(j)}_{2M}\right)^{T}\in\mathbb{C}^{2M}$, $j=1,\ldots,M$ by their components
@@
$m=1,\ldots,2M$. Then the first order approximation of the eigenvalue perturbation $z_{j}^{(1)}$ in (3.12) can be given by
$z_{j}^{(1)}=(\boldsymbol{\mathcal{S}}^{(j)}_{2M})^{T}\,\boldsymbol{\delta}_{2M},\quad j=1,\ldots,M.$ (3.13)
The component $S^{(j)}_{m}$ shows the influence of the perturbation of the $m$th measurement (in this case it is the function value $f(m-1)$) on the $j$th eigenvalue $z_{j}$ of the Hankel matrix pencil (2.6). Using (3.5), we rewrite components $S^{(j)}_{m}$ in the following way
$S^{(j)}_{m}=$ $\frac{1}{\gamma_{j}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M\}}^{\min\{M,m-1\}}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\bigg{)}_{k,j}\,\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\bigg{)}_{m-k,j}$
@@
where $\boldsymbol{z}^{(1)}:=(z_{1}^{(1)},\ldots,z_{M}^{(1)})^{T}$ and $\mathbf{\mathcal{S}}_{M,2M}:=(S_{j,m})_{j,m=1}^{M,2M}\in\mathbb{C}^{M\times 2M}$ be a matrix with the entries $S_{j,m}:=S^{(j)}_{m}$ that quantify the perturbation of the eigenvalues of matrix pencil (2.6).
Finally, we define a structured sensitivity of the eigenvalue $z_{j}$ of matrix pencil (2.6) with respect to the $m$th measurement by
$\eta_{j,m}:=|S_{j,m}|$
for $j=1,\ldots,M$ and $m=1,\ldots,2M$. Then the eigenvalue sensitivity $\eta_{j}$ for the eigenvalue $z_{j}$ with respect to structured perturbations of Hankel matrix pencil (2.6) is defined as the $\ell_{2}$-norm of the vector $\boldsymbol{\mathcal{S}}^{(j)}_{2M}\in\mathbb{C}^{2M}$. Thus we further introduce a vector $\boldsymbol{\eta}:=(\eta_{1},\ldots,\eta_{M})^{T}\in\mathbb{R}^{M}$ with its entries
@@
The ill-disposedness of the eigenvalue $z_{j}$ with respect to structured perturbations can be measured by checking whether $\eta_{j}$ is "large". The entry $\eta_{j}$ determines also the standard deviation of the eigenvalue perturbation $z_{j}^{(1)}$ (see [24] for details).
@@ 4 Recovery of rational functions via Hankel matrix pencil method and sensitivities of the poles
In this section, we describe the main ideas of our new method for the recovery of a rational function of the form (1.1) and present the sensitivity analysis for its poles.
@@ 4.1 Description of the method
Let $r(t)$ be a rational function defined by (1.1) and by $\mathbb{D}$ we denote the unit circle. Following approach in [21], for $k\in\mathbb{Z}$ we consider the integral
@@
which is closely related to the classical Fourier coefficient of a function $r(\mathrm{e}^{\mathrm{i}\varphi})$,
@@
$\widehat{r}_{k}=\begin{cases}\sum\limits_{|z_{j}|<1}\gamma_{j}z_{j}^{-(k+1)},&k\leq-1,\\ -\sum\limits_{|z_{j}|>1}\gamma_{j}z_{j}^{-(k+1)},&k\geq 1.\end{cases}$ (4.2)
From (4.2), we conclude that Fourier coefficients $\widehat{r}_{k}$ with negative indeces $k=-1,-2,\ldots$ contain information about poles inside the unite circle, and Fourier coefficients $\widehat{r}_{k}$ with positive indises $k=1,2,\ldots$ contain information about poles outside the unite circle.
@@
$f(t)=\begin{cases}\sum\limits_{|z_{j}|<1}\gamma_{j}z_{j}^{t},&t\geq 0,\\ -\sum\limits_{|z_{j}|>1}\gamma_{j}z_{j}^{t},&t\leq-2,\end{cases}$ (4.3)
@@
Therefore the rational reconstruction problem (1.1) can be reformulated as the exponential interpolation problem (4.4). Our further goal is to reconstruct an exponential sum (4.3), using samples $f(m)$, $m\in\mathbb{Z}\setminus\{-1\}$ as input data. From (4.3), we get that values $f(m)$ with $m=0,1,2,\ldots$ contain information about nodes $z_{j}$ inside $\mathbb{D}$, and values $f(m)$ with $m=-2,-3,\ldots$ contain information about nodes $z_{j}$ outside $\mathbb{D}$.
Let $M_{1}$ and $M_{2}$ be numbers of nodes $z_{j}$ of the exponential sum (4.3) located inside $\mathbb{D}$ and outside $\mathbb{D}$ respectively. We assume that a number of poles $M$ as well as $M_{1}$ and $M_{2}$ are also unknown. Let $L,N\in\mathbb{N}$ and $M_{1},M_{2}\leq L\leq N$, where $L$ is the upper bound for $M$. We use values $f(k)$, $k=0,\ldots,2N-1$ ($\widehat{r}_{-k}$, $k=1,\ldots,2N$) to reconstruct nodes $z_{j}$ inside $\mathbb{D}$, and values $f(-k)$, $k=2,\ldots,2N+1$ ($\widehat{r}_{k}$, $k=1,\ldots,2N$) to reconstruct nodes $z_{j}$ outside $\mathbb{D}$. Further we rewrite an exponential sum (4.3) in the following way
$f(t)=\begin{cases}\sum\limits_{n=1}^{M_{1}}\gamma^{(-)}_{j_{n}}(z^{(-)}_{j_{n}})^{t},&t\geq 0,\\ -\sum\limits_{n=1}^{M_{2}}\gamma^{(+)}_{j_{n}}(z^{(+)}_{j_{n}})^{t}&t\leq-2,\end{cases}$ (4.5)
i.e. by $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$ we denote poles located inside $\mathbb{D}$ and let $\gamma^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$ be the corresponding coefficients and by $z^{(+)}_{j_{n}}$, $n=1,\ldots,M_{2}$ we denote poles located outside $\mathbb{D}$ and let $\gamma^{(+)}_{j_{n}}$, $n=1,\ldots,M_{2}$ be the corresponding coefficients.
To reconstruct exponential sums (4.5), we use the ESPRIT method described in Section 2. We present some modification of this method for our special parameter estimation problem (4.5). Note that the ESPRIT is not the only method to apply in the frequency domain. We can use the other methods for exponential recovery, for example, MPM (matrix-pencil method) [17] or approximate Prony method [18]. In [21], the author uses the AAK theory [16] in the frequency domain to reconstruct (1.1) and a matrix-valued version of (1.1).
Let $\boldsymbol{z}^{(-)}\in\mathbb{C}^{M_{1}}$ and $\boldsymbol{z}^{(+)}\in\mathbb{C}^{M_{2}}$ be two vectors defined by
$\boldsymbol{z}^{(-)}:=(z^{(-)}_{j_{1}},\ldots,z^{(-)}_{j_{M_{1}}}),\ \ \boldsymbol{z}^{(+)}:=\left(\frac{1}{z_{j_{1}}^{(+)}},\ldots,\frac{1}{z^{(+)}_{j_{M_{2}}}}\right).$
Let us define also diagonal matrices
$\mathbf{\Gamma}^{(-)}_{M_{1}}:=\mathrm{diag}(\gamma^{(-)}_{j_{n}})_{n=1}^{M_{1}},\ \mathbf{\Gamma}^{(+)}_{M_{2}}:=\mathrm{diag}(\gamma^{(+)}_{j_{n}})_{n=1}^{M_{2}},\ \mathbf{Z}_{M_{1}}^{(+)}:=\mathrm{diag}\left(z^{(-)}_{j_{n}}\right)_{n=1}^{M_{1}},\ \mathbf{Z}_{M_{2}}^{(+)}:=\mathrm{diag}\left(\frac{1}{z^{(+)}_{j_{n}}}\right)_{n=1}^{M_{2}}.$
According to the ESPRIT method described in Section 2, to reconstruct nodes $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$, located inside the unit circle, we use the Hankel matrix
@@
and to reconstruct nodes $z^{(+)}_{j_{n}}$, $n=1,\ldots,M_{2}$, located outside the unit circle, we use the Hankel matrix
$\mathbf{H}^{(+)}_{2N-L,L+1}:=(f(-(\ell+k+2)))_{k,\ell=0}^{2N-L-1,L}=\left(\widehat{r}_{k+\ell}\right)_{k=0,\ell=1}^{2N-L-1,L+1}.$
Consider factorizations of these matrices
@@
Since all Vandermonde matrices in (4.6) and (4.7) have full column ranks, the matrix $\mathbf{H}_{2N-L,L+1}^{(-)}$ has rank $M_{1}$ and $\mathbf{H}_{2N-L,L+1}^{(+)}$ has rank $M_{2}$. Therefore in case of exact given data, the numbers $M_{1}$ and $M_{2}$ are reconstructed as numerical ranks of matrices $\mathbf{H}_{2N-L,L+1}^{(-)}$ and $\mathbf{H}_{2N-L,L+1}^{(+)}$ respectively.
First we consider the reconstruction of nodes $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$, located inside the $\mathbb{D}$. Since in this case, the Hankel matrices $\mathbf{H}^{(-)}_{2N-L,L+1}$ have structure similar to the structure of the Hankel matrices $\mathbf{H}_{2N-L,L+1}$ in the Section 2, following ideas of the Section 2, we get that nodes $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$ are reconstructed as eigenvalues of the Hankel matrix pencil
@@
Consider now the case of nodes $z^{(+)}_{j_{n}}$, $n=1,\ldots,M_{2}$, located outside the $\mathbb{D}$. For $s=0,1$, define matrices
@@
and consider their factorizations
@@
Since according to (4.9) and (4.10) we have
$z\,{\mathbf{H}}^{(+)}_{2N-L,L}(0)-{\mathbf{H}}^{(+)}_{2N-L,L}(1)=-\mathbf{V}_{2N-L,M_{2}}(\boldsymbol{z}^{(+)})\,\mathrm{diag}\left(z-\frac{1}{z^{(+)}_{j_{n}}}\right)_{n=1}^{M_{2}}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\mathbf{V}^{T}_{L,M_{2}}(\boldsymbol{z}^{(+)}),$
then inverse $\frac{1}{z^{(+)}_{j_{n}}}$ of nodes $z^{(+)}_{j_{n}}$ for $n=1,\ldots,M_{2}$ are reconstructed as eigenvalues of the Hankel matrix pencil
@@
If we assume that $M_{1}$ and $M_{2}$ are known, we can reconstruct nodes $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$ as eigenvalues of the Hankel matrix pencil ($N=L=M_{1}$ in (4.8))
@@
and inverse $\frac{1}{z^{(+)}_{j_{n}}}$ of nodes $z^{(+)}_{j_{n}}$ for $n=1,\ldots,M_{2}$ are reconstructed as eigenvalues of the Hankel matrix pencil ($N=L=M_{2}$ in (4.11))
@@
Having nodes $z^{(-)}_{j_{n}}$ and $z^{(+)}_{j_{n}}$, coefficients $\gamma^{(-)}_{j_{n}}$ and $\gamma^{(+)}_{j_{n}}$ can be computed as least squares solutions of the overdetermined systems of linear equations
$\sum\limits_{n=1}^{M_{1}}\gamma^{(-)}_{j_{n}}(z^{(-)}_{j_{n}})^{-(k+1)}=\widehat{r}_{-k},\quad k=1,\ldots,2N,$
@@
$-\sum\limits_{n=1}^{M_{2}}\gamma^{(+)}_{j_{n}}(z^{(+)}_{j_{n}})^{-(k+1)}=\widehat{r}_{k},\quad k=1,\ldots,2N,$
@@
1. Note that coefficients $\hat{r}_{k}$ can be also computed applying the quadrature formulas on some grid of points. In that case, we will use the function values of $r$ instead of its Fourier coefficients as input data for the algorithm. Let us chose points $t_{n}=\mathrm{e}^{\frac{2\pi\mathrm{i}n}{4N}}$, $n=0,...,4N-1$ and let $(R_{k})_{k=0}^{4N-1}$ be the DFT of $(r(t_{n}))_{n=0}^{4N-1}$. Then $\hat{r}_{k}$ in (4.1) can be computed as $\hat{r}_{k}=R_{k+2N}$, $k=-2N,...,2N-1$. This question was discussed in details in [21], where the author consider not only the points on the unit circle, but also the Matsubara model as a set of grid points for the evaluation of the transform $\widehat{r}_{k}$. On the one hand, we are interested in the presentation of a modification of the method from [21] and in the explanation of the results of this method for the recovery from noisy data with the use of the sensitivities. In that case, we have to employ the given noisy coefficients $\widehat{r}_{k}$ as input data. On the other hand, we also want to compare the performance of known methods for rational approximation with our new approach based on Hankel pencil. To guarantee fair comparison, we assume to have given noisy values $(r(t_{n}))_{n=0}^{4N-1}$.
2. The method based on using the given coefficients $\hat{r}_{k}$, $k=-2N,...,2N$ has the overall computational cost of $\mathcal{O}(NL^{2})$, which is determined by computation of SVD of Hankel matrices $\mathbf{H}^{(-)}_{2N-L,L+1}$ and $\mathbf{H}^{(+)}_{2N-L,L+1}$. The complexity of the modified algorithm based on using the values $(r(t_{n}))_{n=0}^{4N-1}$ remains the same, since the additional operation computation of FFT of $(r(t_{n}))_{n=0}^{4N-1}$ requires only $\mathcal{O}(N\log N)$ flops.

### v3

@@ 3.1 Sensitivities of Hankel pencil eigenvalues with unstructured perturbations
3.2 Sensitivities of Hankel pencil eigenvalues with structured perturbations
In this subsection, we consider structured perturbations of the GEP (3.1) for the matrix pencil (2.6). Now the perturbation matrices $\Delta_{\mathbf{H}_{M}(0)}$ and $\Delta_{\mathbf{H}_{M}(1)}$ in (3.2) have a particular structure, they are again Hankel matrices. We give a short description of the results from [25]. Let $\boldsymbol{\delta}_{2M}:=(\delta_{0},\delta_{1},\ldots,\delta_{2M-1})^{T}$ be a vector of perturbations and let
$\Delta_{\mathbf{H}_{M}(0)}=(\delta_{k+\ell})_{k,\ell=0}^{M-1,M-1},\ \Delta_{\mathbf{H}_{M}(1)}=(\delta_{k+\ell+1})_{k,\ell=0}^{M-1,M-1}.$
@@ 3.2 Sensitivities of Hankel pencil eigenvalues with structured perturbations
and taking into account (3.14), we define vectors $\boldsymbol{\mathcal{S}}^{(j)}_{2M}:=\left(S^{(j)}_{1},\ldots,S^{(j)}_{2M}\right)^{T}\in\mathbb{C}^{2M}$, $j=1,\ldots,M$ by their components
@@
for $m=1,\ldots,2M$. Then $z_{j}^{(1)}$ as in (3.14) can be presented by
$z_{j}^{(1)}=(\boldsymbol{\mathcal{S}}^{(j)}_{2M})^{T}\,\boldsymbol{\delta}_{2M}.$ (3.15)
The component $S^{(j)}_{m}$ shows the influence of the perturbations of the $m$th measurement (in this case it is the function value f(m1)f(m-1)) on the $j$th eigenvalue $z_{j}$ of the Hankel matrix pencil (2.6). Using (3.6), we rewrite components $S^{(j)}_{m}$ as follows
$S^{(j)}_{m}$= $\frac{1}{\gamma_{j}}\bigg(\,\sum\limits_{k=\max\{1,m-M\}}^{\min\{M,m-1\}}\bigg({\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}}\bigg)_{k,j}\,\bigg(\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\bigg)_{m-k,j}$
@@
where $\boldsymbol{z}^{(1)}:=(z_{1}^{(1)},\ldots,z_{M}^{(1)})^{T}$ and $\mathbf{\mathcal{S}}_{M,2M}:=(S_{j,m})_{j,m=1}^{M,2M}\in\mathbb{C}^{M\times 2M}$ be a matrix with the entries $S_{j,m}:=S^{(j)}_{m}$, which quantify the perturbation of the eigenvalues of the matrix pencil (2.6).
Finally, we define the sensitivity $\eta_{j,m}$ of the eigenvalue $z_{j}$ of the matrix pencil (2.6) with respect to the $m$th measurement by
$\eta_{j,m}:=|S_{j,m}|,$
for $j=1,\ldots,M$ and $m=1,\ldots,2M$. By $\boldsymbol{\eta}_{2M}^{(j)}:=(\eta_{j,1},\ldots,\eta_{j,2M})^{T}\in\mathbb{R}^{2M}$ we denote the corresponding vector of sensitivities for the eigenvalue $z_{j}$. Then the sensitivity $\eta_{j}$ with respect to structured perturbations for the eigenvalue $z_{j}$ of the Hankel matrix pencil (2.6) is defined as the $\ell_{2}$-norm of the vector $\boldsymbol{\mathcal{S}}^{(j)}_{2M}\in\mathbb{C}^{2M}$. Thus, we further define a vector $\boldsymbol{\eta}:=(\eta_{1},\ldots,\eta_{M})^{T}\in\mathbb{R}^{M}$ by its components
@@
The ill-disposedness of the eigenvalue $z_{j}$ with respect to structured perturbations can be measured by checking whether $\eta_{j}$ is ”large“. The entry $\eta_{j}$ determines also the standard deviation of the eigenvalue perturbation $z_{j}^{(1)}$ (see [25] for details).
@@ 4 Recovery of rational functions via Hankel matrix pencil method and sensitivities of the poles
In this section, we describe the main ideas of our new method for the recovery of rational functions (1.1) and present the sensitivity analysis for the poles.
@@ 4.1 Description of the method
Let $r(t)$ be a rational function defined by (1.1) and by $\mathbb{D}:=\{z\in\mathbb{C}:\,|z|=1\}$ we denote the unit circle in the complex plane. Following approach in [22], for $k\in\mathbb{Z}\setminus\{0\}$ we consider the integral
@@
which is closely related to the classical Fourier coefficient $\widehat{r}_{k}$ of the function $r(\mathrm{e}^{\mathrm{i}\varphi})$,
@@
$\widehat{r}_{k}=\begin{cases}\sum\limits_{|z_{j}|<1}\gamma_{j}z_{j}^{-(k+1)},&k=-1,-2,\ldots,\\ -\sum\limits_{|z_{j}|>1}\gamma_{j}z_{j}^{-(k+1)},&k=1,2,\ldots.\end{cases}$ (4.2)
From (4.2), we conclude that Fourier coefficients $\widehat{r}_{k}$ with negative indices $k=-1,-2,\ldots$ contain information about poles located inside the unite circle, and Fourier coefficients $\widehat{r}_{k}$ with positive indices $k=1,2,\ldots$ contain information about poles located outside the unite circle.
@@
$f(t)=\begin{cases}\sum\limits_{|z_{j}|<1}\gamma_{j}z_{j}^{t},&t\geq 0,\\ -\sum\limits_{|z_{j}|>1}\gamma_{j}z_{j}^{t},&t<0,\end{cases}$ (4.3)
@@
Thus, the rational reconstruction problem (1.1) can be reformulated as the exponential interpolation problem (4.4). Our further goal is to reconstruct the exponential sum (4.3), using samples $f(m)$, $m\in\mathbb{Z}\setminus\{-1\}$ as inputs. From (4.3) and (4.4), we conclude that values $f(m)$ with $m=0,1,\ldots$ contain information about nodes $z_{j}$ located inside $\mathbb{D}$, and values $f(m)$ with $m=-2,-3,\ldots$ contain information about nodes $z_{j}$ located outside $\mathbb{D}$.
Let $M_{1}$ and $M_{2}$ be the numbers of nodes $z_{j}$ of $f$ as in (4.3) located inside $\mathbb{D}$ and outside $\mathbb{D}$, respectively. Thus, we can rewrite the exponential sum (4.3) as follows
$f(t)=\begin{cases}\sum\limits_{n=1}^{M_{1}}\gamma^{(-)}_{n}(z^{(-)}_{n})^{t},&t\geq 0,\\ -\sum\limits_{n=1}^{M_{2}}\gamma^{(+)}_{n}(z^{(+)}_{n})^{t}&t<0,\end{cases}$ (4.5)
i.e. by $z^{(-)}_{n}$, $n=1,\ldots,M_{1}$, and $z^{(+)}_{n}$, $n=1,\ldots,M_{2}$, we denote nodes located inside $\mathbb{D}$ and outside $\mathbb{D}$, respectively, and let $\gamma^{(-)}_{n}$, $n=1,\ldots,M_{1}$, and $\gamma^{(+)}_{n}$, $n=1,\ldots,M_{2}$, be the corresponding coefficients.
Next, we present some modification of the ESPRIT method to solve our special parameter estimation problem (4.5). Note that ESPRIT is not the only method to apply in the frequency domain. We can use also the other methods for the exponential recovery, for example, MPM (matrix-pencil method) [18], approximate Prony method [19], or AAK theory [17] (as in [22]).
By $\boldsymbol{z}^{(-)}\in\mathbb{C}^{M_{1}}$ and $\boldsymbol{z}^{(+)}\in\mathbb{C}^{M_{2}}$ we denote two vectors
$\boldsymbol{z}^{(-)}:=(z^{(-)}_{1},\ldots,z^{(-)}_{M_{1}})^{T},\ \ \boldsymbol{z}^{(+)}:=\left(\frac{1}{z_{1}^{(+)}},\ldots,\frac{1}{z^{(+)}_{M_{2}}}\right)^{T},$
and let us also define the following diagonal matrices
$\mathbf{\Gamma}^{(-)}_{M_{1}}:=\mathrm{diag}(\gamma^{(-)}_{n})_{n=1}^{M_{1}},\ \mathbf{\Gamma}^{(+)}_{M_{2}}:=\mathrm{diag}(\gamma^{(+)}_{n})_{n=1}^{M_{2}},\ \mathbf{Z}_{M_{1}}^{(+)}:=\mathrm{diag}\left(z^{(-)}_{n}\right)_{n=1}^{M_{1}},\ \mathbf{Z}_{M_{2}}^{(+)}:=\mathrm{diag}\left(\frac{1}{z^{(+)}_{n}}\right)_{n=1}^{M_{2}}.$
Let further $L,N\in\mathbb{N}$ be given such that $M_{1},M_{2}\leq L\leq N$, and $L$ is the upper bound for $M$. We use values $f(k)$, $k=0,\ldots,2N-1$ ($\widehat{r}_{-k}$, $k=1,\ldots,2N$) for recovering nodes $z^{(-)}_{n}$, $n=1,\ldots,M_{1}$ located inside $\mathbb{D}$, and values f(k)f(-k), $k=2,\ldots,2N+1$ ($\widehat{r}_{k}$, $k=1,\ldots,2N$) for recovering nodes $z^{(+)}_{n}$, $n=1,\ldots,M_{2}$ located outside $\mathbb{D}$. According to the ESPRIT method described in Section 2, to reconstruct nodes $z^{(-)}_{n}$, $n=1,\ldots,M_{1}$, and $z^{(+)}_{n}$, $n=1,\ldots,M_{2}$, we employ the Hankel matrices
@@
and
$\mathbf{H}^{(+)}_{2N-L,L+1}:=(f(-(\ell+k+2)))_{k,\ell=0}^{2N-L-1,L}=\left(\widehat{r}_{k+\ell}\right)_{k=0,\ell=1}^{2N-L-1,L+1},$
respectively. We consider factorizations of these matrices
@@
Since all Vandermonde matrices in (4.6) and (4.7) have full column ranks, $\mathbf{H}_{2N-L,L+1}^{(-)}$ has rank $M_{1}$ and $\mathbf{H}_{2N-L,L+1}^{(+)}$ has rank $M_{2}$. Therefore, in case of exact given data, the numbers $M_{1}$ and $M_{2}$ are reconstructed as numerical ranks of matrices $\mathbf{H}_{2N-L,L+1}^{(-)}$ and $\mathbf{H}_{2N-L,L+1}^{(+)}$, respectively.
Let us start with the case of nodes $z^{(-)}_{n}$, $n=1,\ldots,M_{1}$, located inside $\mathbb{D}$. Since the Hankel matrix $\mathbf{H}^{(-)}_{2N-L,L+1}$ have the structure similar to those of the Hankel matrix $\mathbf{H}_{2N-L,L+1}$ as in (2.2), following ideas of Section 2, we conclude that nodes $z^{(-)}_{n}$, $n=1,\ldots,M_{1}$ can be computed as eigenvalues of the Hankel matrix pencil
@@
Consider now the case of nodes $z^{(+)}_{n}$, $n=1,\ldots,M_{2}$, located outside $\mathbb{D}$. For $s=0,1$, define matrices
@@
and write their factorizations
@@
According to (4.9) and (4.10) we have
$z\,{\mathbf{H}}^{(+)}_{2N-L,L}(0)-{\mathbf{H}}^{(+)}_{2N-L,L}(1)=-\mathbf{V}_{2N-L,M_{2}}(\boldsymbol{z}^{(+)})\,\mathrm{diag}\left(z-\frac{1}{z^{(+)}_{n}}\right)_{n=1}^{M_{2}}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\mathbf{V}^{T}_{L,M_{2}}(\boldsymbol{z}^{(+)}),$
Thus, the inverse $\frac{1}{z^{(+)}_{n}}$ of nodes $z^{(+)}_{n}$ for $n=1,\ldots,M_{2}$ are computed as eigenvalues of the Hankel matrix pencil
@@
If we assume that $M_{1}$ and $M_{2}$ are known, we can reconstruct nodes $z^{(-)}_{n}$, $n=1,\ldots,M_{1}$ as eigenvalues of the Hankel matrix pencil ($N=L=M_{1}$ in (4.8))
@@
and we can reconstruct the inverse $\frac{1}{z^{(+)}_{n}}$ of nodes $z^{(+)}_{n}$, $n=1,\ldots,M_{2}$, as eigenvalues of the Hankel matrix pencil ($N=L=M_{2}$ in (4.11))
@@
Having nodes $z^{(-)}_{n}$ and $z^{(+)}_{n}$, coefficients $\gamma^{(-)}_{n}$ and $\gamma^{(+)}_{n}$ can be computed as least squares solutions of the overdetermined systems of linear equations
$\sum\limits_{n=1}^{M_{1}}\gamma^{(-)}_{n}(z^{(-)}_{n})^{-(k+1)}=\widehat{r}_{-k},\quad k=1,\ldots,2N,$
@@
$\sum\limits_{n=1}^{M_{2}}\gamma^{(+)}_{n}(z^{(+)}_{n})^{-(k+1)}=\widehat{r}_{k},\quad k=1,\ldots,2N,$
@@
1. Note that Fourier coefficients $\hat{r}_{k}$ can also be computed applying quadrature formulas on some grid of points. In that case, we use the function values of $r$ instead of its Fourier coefficients as inputs for the algorithm. Let us chose points $t_{n}=\mathrm{e}^{\frac{2\pi\mathrm{i}n}{4N}}$, n=0,,4N1n=0,...,4N-1 and let $(R_{k})_{k=0}^{4N-1}$ be the DFT of $(r(t_{n}))_{n=0}^{4N-1}$. Then $\hat{r}_{k}$ in (4.1) can be determined by $\hat{r}_{k}=R_{k+2N}$ for k=2N,,2N1k=-2N,...,2N-1. This question was discussed in details in [22], where the author considered not only the points on the unit circle, but also the Matsubara model as a set of grid points for the evaluation of the transform $\widehat{r}_{k}$. On the one hand, we are interested in the presentation of a modification of the method from [22] and in the explanation of the results of this method for the recovery from noisy data with the use of the sensitivities. In that case, we have to employ the given noisy coefficients $\widehat{r}_{k}$ as input data. On the other hand, we also want to compare the performance of known methods for rational approximation with our new approach based on Hankel pencil. To guarantee fair comparison, we assume to have given noisy values $(r(t_{n}))_{n=0}^{4N-1}$.
2. The method based on using the given coefficients $\hat{r}_{k}$, k=2N,,2Nk=-2N,...,2N has the overall computational cost of $\mathcal{O}(NL^{2})$ flops, which is determined by the complexity of the computation of SVDs of Hankel matrices $\mathbf{H}^{(-)}_{2N-L,L+1}$ and $\mathbf{H}^{(+)}_{2N-L,L+1}$. The complexity of the modified algorithm based on employing the values $(r(t_{n}))_{n=0}^{4N-1}$ is $\mathcal{O}(N(L^{2}+\log N))$, since the additional step of the computation of FFT of $(r(t_{n}))_{n=0}^{4N-1}$ requires $\mathcal{O}(N\log N)$ operations.